OpenStudy (mw647):

Please help! Will give medal and fan! What is the concentration of lead ions and sulfide ions in a saturated solution of lead sulfide (PbS) solution at 25ºC (Ksp= 3.0 x 10 -28)?

OpenStudy (jfraser):

what's the equation for the dissociation of lead (II) sulfide, and what's the general format for a KSP equation?

OpenStudy (cuanchi):

https://answers.yahoo.com/question/index?qid=20090319121817AAxdnkh Dr.A answered 6 years ago Let x = mol/L of PbS that dissolve. This will give x mol/L of Pb2+ and x mol/L of S2- Ksp = [Pb2+][S2-] = (x)(x) = x^2 x = molar solubility = square root Ksp=1.74 x 10^-14 M that is the concentration of each ion in the saturated solution